OpenStudy (anonymous):

A bottle of cola has 39 grams of sugar. A bottle of diet cola contains no sugar. If you drink two bottles of regular cola and one bottle of diet cola, how many grams of sugar will you consume?

OpenStudy (nurali):

2 bottle of regular = 2 * 39grams 1 bottle of diet = 1 * 0 grams 2*39g + 1*0g 78g + 0g 78g

OpenStudy (anonymous):

wow that was fast thank you Nurali

OpenStudy (nurali):

Anytime.